OpenStudy (anonymous):

A person purchases a television set for Rs 16000. It's estimated life is 20 years. If yearly depreciation is assumed to be constant. Find the rate of depreciation and price after 8 years.

OpenStudy (amistre64):

the depreciation is constant sooo; 16000/20

OpenStudy (amistre64):

since 8 years of depreciation have accumulated .... the price is 16000 - 8(dep)

OpenStudy (anonymous):

t8= 16000+(8-1)(-800) = 16000-5600 = 10400

OpenStudy (anonymous):

but answer given is 9600

OpenStudy (amistre64):

why would you only do it for 7 years?

OpenStudy (anonymous):

isn't this an arithmetic sequence with first term 16000 and common difference - 800 ?? so the formula of t8 would be t8 = a+(n-1)d ?

OpenStudy (amistre64):

if you insist on starting at t1, then t(1+8) = t9

OpenStudy (amistre64):

if you start at t0, then there is no need to subtract 1

OpenStudy (amistre64):

t0 = 16000 - 800(0) t1 = 16000 - 800(1) t2 = 16000 - 800(2) ... t8 = 16000 - 800(8)

OpenStudy (anonymous):

:) got u ... thanks !!

OpenStudy (amistre64):

youre welcome